OpenStudy (anonymous):

solve the equation. 1/ 3x + 9 - 2/ x + 3 = 2

OpenStudy (anonymous):

\[\frac{ 1 }{ 3x + 9 } - \frac{ 2 }{ x + 3 } = 2\]

OpenStudy (jhannybean):

oh okay.

OpenStudy (anonymous):

?

OpenStudy (luigi0210):

you got this?

OpenStudy (anonymous):

No. I don't know why she said oh okay.

OpenStudy (bahrom7893):

I GOT THIS YALL!

OpenStudy (jhannybean):

No lol. im guessing we're going to multiply the fractions with the denominators of both... can't find an easy common denominator :(

OpenStudy (luigi0210):

Alright,YEA BAHROM!

OpenStudy (luigi0210):

here I'll start you guys off: